# Teams of LLM Agents can Exploit Zero-Day Vulnerabilities

### v1

@@ 1 Introduction
AI agents are rapidly becoming more capable. They can now solve tasks as complex as resolving real-world GitHub issues [1] and real-world email organization tasks [2]. However, as their capabilities for benign applications improve, so does their potential in dual-use settings.
Of the dual-use applications, hacking is one of the largest concerns [3]. As such, recent work has explored the ability of AI agents to exploit cybersecurity vulnerabilities [4, 5]. This work has shown that simple AI agents can autonomously hack mock “capture-the-flag” style websites and can hack real-world vulnerabilities when given the vulnerability description. However, they largely fail when the vulnerability description is excluded, which is the zero-day exploit setting [5]. This raises a natural question: can more complex AI agents exploit real-world zero-day vulnerabilities?
In this work, we answer this question in the affirmative, showing that teams of AI agents can exploit real-world zero-day vulnerabilities. To show this, we develop a novel multi-agent framework for cybersecurity exploits, extending prior work in the multi-agent setting [6, 7, 8]. We call our technique HPTSA, which (to our knowledge) is the first multi-agent system to successfully accomplish meaningful cybersecurity exploits.
@@ 2.1 Computer Security
Computer security is a broad field. In this work, we focus specifically on vulnerability exploitation, which is just one part of the wider field of computer security and even attacks. For example, after a vulnerability is exploited, an attacker must typically perform lateral movement to cause harm [9].
In this work, we focus on vulnerabilities in a computer system that are unknown to the deployer of the computer system. Unfortunately, the term of these vulnerabilities vary from source to source, but we refer to these vulnerabilities as zero-day vulnerabilities (0DV). This is in contrast to one-day vulnerabilities (1DV), where the vulnerability is disclosed but unpatched.
Zero-day vulnerabilities are particularly harmful because the system deployer cannot proactively put mitigations in place against these vulnerabilities [10]. We focus specifically on web vulnerabilities in this work, which are often the first attack surface into more in depth attacks [11].
One important distinction within vulnerabilities is the class of vulnerability and the specific instance of the vulnerability. For example, server-side request forgery (SSRF) has been known as a class of vulnerability since at least 2011 [12]. However, one of the biggest hacks of all time that occurred in 2021 (10 years after) hacked Microsoft, now a multi-trillion dollar company that invests about a billion dollars a year in computer security [13], used an SSRF [14].
@@ 2.2 AI Agents and Cybersecurity
AI agents have become increasingly powerful and can perform tasks as complex as solving real-world GitHub issues [1]. In this work, we focus on AI agents solving complex, real-world tasks. These agents are now almost exclusively powered by tool-enabled LLMs [15, 16]. The basic architecture of these agents involves an LLM that is given a task and carries out that task by using tools via APIs. We provide a more detailed overview of AI agents in Section 8.
Recent work has explored AI agents in the context of cybersecurity, showing that they can exploit “capture-the-flag” style vulnerabilities [4] and one-day vulnerabilities when given a description of the vulnerability [5]. These agents work via the simple ReAct-style iteration, where the LLM takes an action, observes the response, and repeats [17].
@@ 3 HPTSA: Hierarchical Planning and Task-Specific Agents
As mentioned, ReAct-style agents iterate by taking actions, observing the response, and repeating. Although successful for many kinds of tasks, the repeated iteration can make long-term planning fail because 1) the context can extend rapidly for cybersecurity tasks, and 2) it can be difficult for the LLM to try many different exploits. For example, prior work has shown that if an LLM agent attempts one type of vulnerability, backtracking to try another type of vulnerability is challenging for a single agent [5].
@@ 3.1 Overall Architecture
Our first component is the hierarchical planner, which explores the environment (i.e., website). After exploring the environment, it determines the set of instructions to send to the team manager. For example, the hierarchical planner may determine that the login page is susceptible to attacks and focus on that.
Our second component is a team manager for the task-specific agents. It determines which specific agents to use. For example, it may determine that a SQLi expert agent is the appropriate agent to use on a specific page. Beyond choosing which agents to use, it also retrieves the information from previous agent runs. It can use this information to rerun task-specific agents with more detailed instructions or run other agents with information from the previous runs.
@@ 3.2 Task-Specific Agents
In order to increase the performance of teams of agents in the cybersecurity setting, we designed task-specific, expert agents. We designed 6 total expert agents: XSS, SQLi, CSRF, SSTI, ZAP, and a “generic” web hacking agent. Our AI agents have: 1) access to tools, 2) access to documents, and 3) a prompt.
For the tools, all agents had access to Playwright (a browser testing framework to access the websites), the terminal, and file management tools. The ZAP agent also had access to ZAP [18]. The agents accessed the websites via Playwright. We manually ensured that the agents did not search for the vulnerabilities via search engines or otherwise.
Unfortunately, certain tools that may be useful do not work well with the OpenAI assistants so we excluded them. For example, sqlmap, a framework for testing for potential SQL injections, may be useful for the SQLi agent. However, as it runs timing attacks, it does not work with the 10 minute limit the OpenAI assistants have.
@@
Finally, for the prompt, we used the same prompt template but modified them for each vulnerability.
@@ 3.3 Implementation
For our specific implementation for HPTSA for web vulnerabilities, we used the OpenAI assistants API in conjunction with LangChain and LangGraph. We used GPT-4 for all experiments in our work, since prior work has shown that GPT-4 is far more proficient at hacking tasks compared to other models [4, 5].
We used LangGraph’s functionality to create a graph of agents and passed messages between agents using LangGraph. The individual agents were implemented with a conjunction of OpenAI Assistants and LangChain.
@@ 4 Benchmark of Zero-Day Vulnerabilities
Second, we focused on web vulnerabilities that we could reproduce and with a specific trigger. Many non-web vulnerabilities require complex environments to set up or have vague conditions for success. For example, prior work tests vulnerabilities in Python packages that, when included, allow for arbitrary code execution. This is difficult to test, since it requires a testing framework that includes the code. In contrast, the web vulnerabilities had clear pass or fail measures.
Based on these criteria, we collected 15 web vulnerabilities. Our vulnerabilities include many vulnerability types, including XSS, CSRF, SQLi, arbitrary code execution, and others. They are all of severity medium or higher (including high severity and critical vulnerabilities).
Table 1: List of vulnerabilities we consider and their description. ACE stands for arbitrary code execution and RCE stands for remote code execution. Further details are given in Table 2.
@@
flusity-CMS CSRF CSRF vulnerability in flusity-CMS v2.33, allows remote attackers to execute arbitrary code
@@
WPZest SQLi Improper neutralization of special elements used in an SQL Command
@@
Reportico ACE Issue allows attacker to execute arbitrary code and obtain sensitive information via the sessionid
@@
PrestaShop information leakage Random secure_key parameter allows arbitrary users to download any invoice
@@
WPZest SQLi CVE-2024-32135 04/15/2024 7.6 (high)
@@
Reportico ACE CVE-2024-31556 05/14/2024 6.5 (medium)
@@ 5.1 Experimental Setup
Metrics. We measure the success of our agents with the pass at 5 and pass at 1 (i.e., overall success rate). Unlike for many other tasks, if a single attempt is successful, the attacker has successfully exploited the system. Thus, pass at 5 is our primary metric.
In order to determine if the agent successfully exploited a vulnerability, we manually verified the trace to ensure that the requisite set of actions were taken to exploit the vulnerability.
@@
Baselines. In addition to testing our most capable agent, we additionally tested several variants of our agents. As an upper bound on performance, we tested the one-day agent used by Fang et al [5], in which the agent is given the description of the vulnerability. As a lower bound on performance, we tested the one-day agent without the vulnerability description. Finally, we test the open-source vulnerability scanners ZAP [18] and MetaSploit [19]. We further test on several ablations of HPTSA, which we describe below.
For all agents, we used gpt-4-0125-preview (i.e., GPT-4 Turbo) which has training data up to December 2023 (according to OpenAI). Prior work has shown that other models, including GPT-3.5 and high-performing open-source models, perform poorly on cybersecurity exploits [4, 5]. As such, we did not test other models.
Vulnerabilities. We tested all of our agents on the vulnerabilities we collected, described in Table 1. To ensure that no real users were harmed, we reproduced these vulnerabilities in a sandboxed environment. To reiterate, all vulnerabilities were past the GPT-4 cutoff date at the time of experimentation. Furthermore, all of our vulnerabilities were of severity medium or higher, and we benchmarked against a variety of vulnerabilities.
@@ 5.2 End-to-End results
Figure 2: Pass at 5 and overall success rate (pass at 1) for open-source vulnerability scanners, GPT-4 with no description, HPTSA, and GPT-4 w/ desc.
We measured the overall success rate of our highest performing agent (HPTSA), the agent with the vulnerability description (GPT-4 w/ desc.), the agent without the vulnerability description (GPT-4 no desc.), and the open-source vulnerability scanners. We show results in Figure 2.
[…]
@@ 5.3 Ablation studies
Figure 3: Pass at 5 and overall success rate (pass at 1) for HPTSA without documents and without expert agents.
To further understand the capabilities of our GPT-4 agent, we tested two ablations of our agents: 1) when replacing the task-specific agents with a single generic cybersecurity agent, and 2) when removing the documents from the task-specific agents. We show results in Figure 3.
@@ 6.1 Success Case Studies
      Run 1: The agent attempted a SQL injection attack on the login page, which did not work.
@@
      Run 1: The agent successfully logged in and navigated to the menu creation endpoint. There, it took the steps to create a menu on its own. It then verified that a new menu was created, and crafted a CSRF payload that recreates those steps, exploiting CVE-2024-24524.
@@
      Run 3: The agent logged in and navigated to the post creation page, again attempting to craft a payload that would create a new post. However, the payload again did not work.
@@
We also note that the task-specific agents can now focus specifically on the vulnerability and does not need to backtrack, as the backtracking is in the purview of the supervisor agent. Prior work observed that a single agent often gets confused in backtracking [5], which HPTSA resolves.
@@ 6.2 Unsuccessful Case Studies
Our results suggest that our agents could be further improved by forcing the expert agents to work on specific pages and exploring endpoints that are not easily accessible, either by brute force or other techniques.
@@ 7 Cost Analysis
In line with prior work [4, 5], we measure the cost of our GPT-4 agent. Similar to prior work, our estimates are not meant to reflect the end-to-end cost of complete, real-world hacking tasks. We provide these estimates so that the cost of our agents can be put in the context of prior work.
As mentioned, we measure the cost of our agents by tracking the input and output tokens. At the time of writing, GPT-4 costs \$30 per million output tokens and \$10 per million input tokens. Note that we use GPT-4 Turbo with training data up to December 2023.
@@ 8 Related Work
In this work, we focus on AI agents and cybersecurity. The closest works to ours shows that ReAct-style AI agents can hack “capture-the-flag” toy websites and vulnerabilities when given a description of the vulnerability [4, 5]. However, these agents fare poorly in the zero-day setting. In particular, it is challenging for agents to backtrack after exploring a dead end. We show in our work that teams of AI agents can autonomously exploit zero-day vulnerabilities.
@@
The human uplift setting focuses on using AI (typically LLMs) to aid humans in cybersecurity tasks. For example, recent work has shown that LLMs can aid humans in penetration testing and malware generation [20, 21]. This work is especially important in the setting of “script kiddies” who deploy malware without special expertise. Based on this work, and the work on AI agents, researchers have also speculated on the societal implications of AI on cybersecurity [3, 22].
@@
AI agents. AI agents have becoming increasing powerful and popular. Recent highly capable AI agents are largely based on LLMs [17, 16] and can now perform tasks as complex as solving real-world GitHub issues [1]. There have been hundreds of papers on improving AI agents, ranging from prompting techniques [23, 24], planning techniques [25, 26], adding documents and memory [27], domain-specific agents [28], and many more [15]. Particularly related to our work is the field of multi-agent systems [6, 7, 8]. However, to the best of our knowledge, our work is the first to introduce a real-world AI agent system based on hierarchical planning and task-specific agents.
@@
Security of AI agents. A related area of work is the security of AI agents themselves [29, 30, 31, 32, 33, 34]. Deployers of AI agents may want to limit the tasks that the AI agent can do (e.g., restricting the ability to perform cybersecurity attacks) and protect the agent against malicious attackers. Unfortunately, recent work has shown that it is simple to bypass protections in LLMs, such as by fine-tuning away protections [32, 34, 33]. AI agents can also be attacked via indirect prompt injection attacks [35, 36, 37]. This line of work is orthogonal to ours.
9 Conclusions and Limitations

### v2

@@ Abstract
Teams of LLM Agents can Exploit Zero-Day Vulnerabilities
Yuxuan Zhu1, Antony Kellermann2, Akul Gupta1, Philip Li1, Richard Fang1, Rohan Bindu1, Daniel Kang1 University of Illinois Urbana Champaign 1{yxx404, akulg3, philipl2, rrfang2, bindu2, ddkang}@illinois.edu, 2antony@aokellermann.dev
@@ 1 Introduction
AI agents are rapidly becoming more capable. They can now solve tasks as complex as resolving real-world GitHub issues Yang et al. (2024b) and real-world email organization tasks Roth and Davis (2024). However, as their capabilities for benign applications improve, so does their potential in dual-use settings.
Of the dual-use applications, hacking is one of the largest concerns Lohn and Jackson (2022). As such, recent work has explored the ability of AI agents to exploit cybersecurity vulnerabilities Fang et al. (2024b, a). This work has shown that simple AI agents can autonomously hack mock “capture-the-flag” style websites and can hack real-world vulnerabilities when given the vulnerability description. However, they largely fail when the vulnerability description is excluded, which is the zero-day exploit setting Fang et al. (2024a). This raises a natural question: can more complex AI agents exploit real-world zero-day vulnerabilities?
In this work, we answer this question in the affirmative, showing that teams of AI agents can exploit real-world zero-day vulnerabilities. To show this, we develop a novel multi-agent framework for cybersecurity exploits, extending prior work in the multi-agent setting Liu et al. (2023b); Chen et al. (2023); Zhang et al. (2023). We call our technique HPTSA, which (to our knowledge) is the first multi-agent system to successfully accomplish meaningful cybersecurity exploits.
@@ 2.1 Computer Security
In this work, we focus on the vulnerability exploitation of computer systems. A vulnerability in a computer system is flaw in that system that allows behaviors unintended by the creator of the system, typically for malicious use. Exploiting the vulnerability consists of detecting the vulnerability and performing the necessary actions to take advantage of the vulnerability.
We focus on vulnerabilities in a computer system that are unknown to the deployer of the system. Unfortunately, the term of these vulnerabilities vary from source to source, but we refer to these vulnerabilities as zero-day vulnerabilities (0DV). This is in contrast to one-day vulnerabilities (1DV), where the vulnerability is disclosed but unpatched. Namely, a 1DV is known to the attacker.
Zero-day vulnerabilities are particularly harmful because the system deployer cannot proactively put mitigations in place against these vulnerabilities Bilge and Dumitraş (2012). We focus specifically on web vulnerabilities in this work, which are often the first attack surface into more in depth attacks Setiawan and Setiyadi (2018).
One important distinction within vulnerabilities is the class of vulnerability and the specific instance of the vulnerability. For example, server-side request forgery (SSRF) has been known as a class of vulnerability since at least 2011 Fung and Lee (2011). However, one of the biggest hacks of all time that occurred in 2021 (10 years after) hacked Microsoft, now a multi-trillion dollar company that invests about a billion dollars a year in computer security Microsoft (2024), used an SSRF Kost (2023).
@@ 2.2 AI Agents and Cybersecurity
AI agents have become increasingly powerful and can perform tasks as complex as solving real-world GitHub issues Yang et al. (2024b). In this work, we focus on AI agents solving complex, real-world tasks. These agents are now almost exclusively powered by tool-enabled LLMs Parisi et al. (2022); Weng (2023). The basic architecture of these agents involves an LLM that is given a task and carries out that task by using tools via APIs. We provide a more detailed overview of AI agents in Section 8.
Recent work has explored AI agents in the context of cybersecurity, showing that they can exploit “capture-the-flag” style vulnerabilities Fang et al. (2024b); Zhang et al. (2024) and one-day vulnerabilities when given a description of the vulnerability Fang et al. (2024a). These agents work via the ReAct-style iteration, where LLMs take an action, observe the response, and repeat Yao et al. (2022).
@@ 3 HPTSA: Hierarchical Planning and Task-Specific Agents
As mentioned, ReAct-style agents iterate by taking actions, observing the response, and repeating. Although successful for many kinds of tasks, the repeated iteration can make long-term planning for cybersecurity tasks fail because 1) the context can extend rapidly for cybersecurity tasks, and 2) it can be difficult for the LLM to try many different exploits. For example, prior work has shown that if an LLM agent attempts one type of vulnerability, backtracking to try another type of vulnerability is challenging for a single agent Fang et al. (2024a).
@@ 3.1 Overall Architecture
Our first component is the hierarchical planner, which explores the environment (i.e., websites). After exploring the environment, it determines the set of instructions to send to the team manager. For example, the hierarchical planner may determine that the login page is susceptible to attacks and focus on that.
Our second component is a team manager for the task-specific agents. It determines which specific agents to use. For example, it may determine that a SQLi expert agent is the appropriate agent to use on a specific page. Beyond choosing which agents to use, it also retrieves the information from previous agent runs. It can use this information to rerun task-specific agents with more detailed instructions or run other agents.
@@ 3.2 Task-Specific Agents
In order to increase the performance of teams of agents in the cybersecurity setting, we designed task-specific, expert agents. We designed 6 total expert agents: XSS, SQLi, CSRF, SSTI, ZAP, and a “generic” web hacking agent. Our AI agents have: 1) access to tools, 2) access to documents, and 3) specific prompts.
For the tools, all agents had access to Playwright (a browser testing framework to access the websites), the terminal, and file management tools. The ZAP agent also had access to ZAP Bennetts (2013), while the SQLi agent had access to sqlmap sqlmap (2024). The agents accessed the websites via Playwright. We manually ensured that the agents did not search for the vulnerabilities via search engines or otherwise.
@@
Finally, for the prompt, we used the same prompt template. We further customized them for each vulnerability to give agents the necessary information, such as a user account, to execute the attack.
@@ 3.3 Implementation
In our specific implementation for HPTSA for web vulnerabilities, we used the LangChain and LangGraph library in conjunction to APIs of Fireworks and OpenAI assistants. We used LangGraph’s functionality to create a graph of agents and passed messages between agents using LangGraph. The individual agents were implemented with a conjunction of OpenAI Assistants, Fireworks, and LangChain.
@@ 4 Benchmark of Zero-Day Vulnerabilities
flusity-CMS CSRF CSRF vulnerability in flusity-CMS v2.33, allows ACE
@@
Dolibarr SQLi Improper neutralization of special elements used in an SQL Command
@@
Zabbix privilege escalation Improper input sanitization leads to a privilege escalation
@@
PrestaShop information leakage Random secure_key parameter allows any user to download any invoice anonymously
Table 1: List of vulnerabilities we consider and their description. ACE stands for arbitrary code execution. Further details are given in Table 2.
@@
Dolibarr SQLi CVE-2024-5314 05/24/2024 9.1 (critical)
@@
Zabbix privilege escalation CVE-2024-22120 05/14/2024 9.1 (critical)
@@
Second, we focused on web vulnerabilities with a specific trigger. Many non-web vulnerabilities require complex environments to set up or have vague conditions for success. For example, prior work tests vulnerabilities in Python packages that, when included, allow for arbitrary code execution. This is difficult to test, since it requires a testing framework that includes the code. In contrast, the web vulnerabilities had clear pass or fail measures.
Finally, we included only vulnerabilities that we can exploit manually to ensure the reproducibility of our benchmark. Some vulnerabilities cannot be replicated if the specific version of the required package is no longer officially available.
Based on these criteria, we collected 14 web vulnerabilities. Our vulnerabilities include many vulnerability types, including XSS, CSRF, SQLi, arbitrary code execution, and others. They are all of severity medium or higher (including high severity and critical vulnerabilities).
@@ 5.1 Experimental Setup
Metrics. Recall that our work focuses on vulnerability exploitation as opposed to detection. Thus, we measure the success of our agents exploiting the vulnerabilities at hand. To measure this, we manually checked the agent traces to confirm that the vulnerabilities were successfully exploited.
We measure the success of our agents with the pass at 5 and pass at 1 (i.e., overall success rate). Unlike for many other tasks, if a single attempt is successful, the attacker has successfully exploited the system. Thus, pass at 5 is our primary metric.
@@
Baselines. In addition to testing our most capable agent, we additionally tested several variants of it.
As an upper bound on performance, we tested the one-day agent used by Fang et al. (2024a), in which the agent is given the description of the vulnerability. This agent has strictly more information than our agent, since it knows the vulnerability. We refer to this agent as 1DV agent.
As a lower bound on performance, we tested the one-day agent without the vulnerability description. Finally, we test the open-source vulnerability scanners ZAP Bennetts (2013) and MetaSploit Kennedy et al. (2011). We further test on several ablations of HPTSA, which we describe below.
Models. For HPTSA, we used both proprietary and open-source models, including
  1. 1.
    gpt-4-0125-preview Achiam et al. (2023)
  2. 2.
    llama-3.1-405B Dubey et al. (2024)
  3. 3.
    qwen 2.5 72B Yang et al. (2024a)
Vulnerabilities. We tested all of our agents on the vulnerabilities we collected, described in Table 1. To ensure that no real users were harmed, we reproduced these vulnerabilities in a sandboxed environment. Furthermore, all of our vulnerabilities were of severity medium or higher, and we benchmarked against a variety of vulnerabilities.
@@ 5.2 End-to-End results
Figure 2: Pass at 5 and overall success rate (pass at 1) for HPTSA with various models.
(a) Pass at 5
(b) Overall success rate (pass at 1)
Figure 3: Pass at 5 and overall success rate (pass at 1) for open-source vulnerability scanners, GPT-4 with no description, HPTSA, and GPT-4 with description.
We measured the overall success rate of our highest performing agent (HPTSA) with different models. We also compared HPTSA with the agent with vulnerability descriptions (1DV agent), the agent without the vulnerability description (GPT-4 no desc.), and the open-source vulnerability scanners.
As shown in Figure 2, HPTSA with GPT-4 reaches the highest success rate, achieving a 42% pass at 5 and an 18% pass at 1. In contrast, open-source models failed to exploit any vulnerability. We observed that open-source models had a higher rate of refusals (e.g., 31% for llama) and often repeatedly attempted the same incorrect approach. As these results show, GPT-4 powered agents can successfully exploit real-world vulnerabilities in the zero-day setting. Our results resolve an open question in prior work, showing that a more complex and structured agent setup (HPTSA) can exploit zero-day vulnerabilities effectively Fang et al. (2024a).
@@ 5.3 Ablation studies
Figure 4: Pass at 5 and overall success rate (pass at 1) for HPTSA without documents, task-specific agents, or hierarchical structure.
To further understand the capabilities of our agents, we tested two ablations of our agents: 1) when replacing the task-specific agents with a single generic cybersecurity agent, 2) when removing the documents from the task-specific agents. We show results in Figure 4, and 3) when using task-specific agent randomly without the hierarchical structure.
@@ 6.1 Success Case Studies
      Run 1: The agent attempted a SQL injection attack on the login page, which failed.
@@
      Run 1: The agent successfully logged in and navigated to the menu creation endpoint. There, it took the steps to create a menu. It then verified that a new menu was created, and crafted a CSRF payload that recreates those steps, exploiting CVE-2024-24524.
@@
      Run 3: The agent logged in and navigated to the post creation page, again attempting to craft a payload that would create a new post. However, the payload did not work.
@@
We also note that the task-specific agents can now focus specifically on the vulnerability and does not need to backtrack, as the backtracking is in the purview of the supervisor agent. Prior work observed that a single agent often gets confused in backtracking Fang et al. (2024a), which is resolved by HPTSA.
@@ 6.2 Unsuccessful Case Studies
Our results suggest that our agents could be further improved by forcing the expert agents to work on specific types of pages and exploring endpoints that are not easily accessible, either by brute force or other techniques.
@@ 7 Cost Analysis
In line with prior work Fang et al. (2024b, a), we measure the cost of our HPTSA. Similar to prior work, our estimates are not meant to reflect the end-to-end cost of complete, real-world hacking tasks. We provide these estimates so that the cost of our agents can be put in the context of prior work.
As mentioned, we measure the cost of our agents by tracking the input and output tokens. At the time of writing, GPT-4 costs \$30 per million output tokens and \$10 per million input tokens. For open-source models, we used Fireworks API, costing \$3 per million tokens for Llama-3.5-405B and \$0.9 per million tokens for Qwen-2.5-72B.
Model Cost / run Cost / success
gpt-4-0125-preview \$4.39 \$24.4
llama-3.1-405B \$0.30 N/A (no success)
qwen-2.5-72B \$1.41 N/A (no success)
Table 3: Average cost per run of HPTSA.
@@ 8 Related Work
In this work, we focus on AI agents and cybersecurity. The closest works to ours shows that ReAct-style AI agents can hack “capture-the-flag” toy websites and vulnerabilities when given a description of the vulnerability Fang et al. (2024b, a). However, these agents fare poorly in the zero-day setting. In particular, it is challenging for agents to backtrack after exploring a dead end. We show in our work that teams of AI agents can autonomously exploit zero-day vulnerabilities. Our findings are of broader relevance to the community, as governmental agencies US (2025); UK (2024), industrial labs Weidinger et al. (2024); Anthropic (2024), and other parties are interested in measuring cybersecurity capabilities of AI agents.
@@
The human uplift setting focuses on using AI (typically LLMs) to aid humans in cybersecurity tasks. For example, recent work has shown that LLMs can aid humans in penetration testing and malware generation Happe and Cito (2023); Hilario et al. (2024). This work is especially important in the setting of “script kiddies” who deploy malware without special expertise. Based on this, and the work on AI agents, researchers have also speculated on societal implications of AI on cybersecurity Lohn and Jackson (2022); Handa et al. (2019).
@@
AI agents. AI agents have becoming increasing powerful and popular. Recent, highly capable AI agents are largely based on LLMs Yao et al. (2022); Weng (2023) and can now perform tasks as complex as solving real-world GitHub issues Yang et al. (2024b). There have been hundreds of papers on improving AI agents, ranging from prompting techniques Wei et al. (2022); Yao et al. (2024), planning techniques Shinn et al. (2024); Liu et al. (2023a), adding documents and memory Nuxoll and Laird (2012), domain-specific agents He et al. (2024), and many more Parisi et al. (2022). The field of multi-agent systems is particularly related to our work Liu et al. (2023b); Chen et al. (2023); Zhang et al. (2023). However, to the best of our knowledge, our work is the first to introduce a real-world AI agent system based on hierarchical planning and task-specific agents.
@@
Security of AI agents. A related area of work is the security of AI agents themselves Greshake et al. (2023a); Kang et al. (2023); Zou et al. (2023); Zhan et al. (2023); Qi et al. (2023); Yang et al. (2023). Deployers of AI agents may want to limit the tasks that the AI agent can do (e.g., restricting the ability to perform cybersecurity attacks) and protect the agent against malicious attackers. Unfortunately, recent work has shown that it is simple to bypass protections in LLMs, such as by fine-tuning away protections Zhan et al. (2023); Yang et al. (2023); Qi et al. (2023). AI agents can also be attacked via indirect prompt injection attacks Greshake et al. (2023b); Yi et al. (2023); Zhan et al. (2024). This line of work is orthogonal to ours.
9 Conclusions